# A Dynamical-Time Framework for the Dynamics of Charged Particles

### v1

@@ Abstract
We present a dynamical framework for modeling the motion of point-like charged particles, with or without mass, in general external electromagnetic fields. A key feature of this formulation is the treatment of the time coordinate as a dynamical variable. The framework applies to the relativistic regime while consistently admitting a nonrelativistic limit. We also present a representation of particle trajectories in velocity space, which provides a clear insight into the nature and asymptotic behavior of the dynamics. As an application, we compare the motion of massive and massless particles in a constant electromagnetic field and find that, for identical field configurations, their asymptotic behavior is independent of both mass and initial conditions. Finally, we explore the computational advantages of the dynamical-time formulation over the conventional uniform-time approach in two study cases: an uniform electromagnetic field and an elliptically polarized wave propagating along a uniform magnetic field. In both scenarios, the proposed scheme exhibits improvements in accuracy and computational efficiency.
@@ I Introduction
A substantial body of literature is devoted to the classical dynamics of electrically charged particles, with or without spin and irrespective of their mass; a comprehensive review lies beyond the scope of the present study. For a partial survey, we refer the reader to Ref. [moralesBehaviourChargedSpinning2018]. For complementary approaches not covered therein, see Refs. [Plyushchay:1988ws, Plyushchay:1988wx, Nesterenko:1991av], as well as the extensive bibliography in Ref. [arreaga-garcia2014equations]. Of particular relevance to the present work are the Lagrangian formulations proposed in Refs. [brinkLocalSupersymmetrySpinning1976, brinkLagrangianFormulationClassical1977, balachandranClassicalDescriptionParticle1977], in which the kinetic term is decoupled from the mass, thereby naturally encompassing the massless case.
In nonlinear dynamics, the behavior of charged particles in electromagnetic fields represents a fruitful area of research [akhiezerDynamicsHighenergyCharged1995, ramDynamicsChargedParticles2010]. For instance, massive charged particles in a dipolar magnetic field exhibit dynamical behavior ranging from periodic and quasiperiodic to chaotic and hyperchaotic in both the low-velocity [dealcantarabonfimChaoticHyperchaoticMotion2000] and relativistic regimes [shebalinStormerRegionsAxisymmetric2004, xiePeriodQuasiperiodChaos2020]. Extending such studies to massless particles could therefore provide valuable insights into ultra-relativistic dynamics. However, a primary difficulty for this extension is that the Lorentz factor diverges when the particle mass vanishes, which challenges the application of standard approaches [petriFullyImplicitNumerical2017].
@@ II Dynamics of charged particles
In this section, we review the classical dynamics of charged particles, mostly based on Refs. [moralesBehaviourChargedSpinning2018, moralesQuantumChargedSpinning2019].
@@
$S_{\mathrm{B}}[x,e]=-\int_{\mathcal{C}}\mathrm{d}\lambda\left(\frac{\dot{x}^{2}}{2e}+\frac{em^{2}}{2}+qA_{\mu}(x)\dot{x}^{\mu}\right),$ (2)
@@
In the action (2), the mass is decoupled from the kinetic term, appearing as a free parameter that multiplies the particle’s worldline length. Consequently, the massless sector is straightforwardly realized. Moreover, one may even substitute $m^{2}\rightarrow-m^{2}$ to describe tachyonic particles. [^1] The action (2) can be interpreted as a gravitational theory in $0+1$ dimensions coupled to the massless scalar fields $x^{\mu}(\lambda)$ [brinkLocalSupersymmetrySpinning1976], where $m^{2}$ plays the role of a cosmological constant that is positive, negative (if $m^{2}\to-m^{2}$), or zero.
@@
The Hamiltonian formulation of Brink’s theory, in its most general form describing a charged spinning particle, was developed in Ref. [moralesQuantumChargedSpinning2019]. In the present case, when spin is neglected, the first-order action takes the form
$S[e,x,P]=\int\mathrm{d}\lambda\left(P_{\mu}\dot{x}^{\mu}-eH(x,P)\right),$ (5)
where $P_{\mu}\equiv\delta S_{\mathrm{B}}/\delta\dot{x}^{\mu}$ are the canonical momenta, and the Hamiltonian is given by
$H(x,P)=\frac{1}{2}\left(m^{2}-\eta^{\mu\nu}\left(P_{\mu}+qA_{\mu}\right)\left(P_{\nu}+qA_{\nu}\right)\right)\,.$ (6)
@@
The Hamilton equations lead to the following dynamical system for the canonically conjugate pair $(x^{\mu},P_{\nu})$:
@@
$\dot{P}_{\mu}$ $=\{P_{\mu},eH\}=qe\eta^{\sigma\rho}\partial_{\mu}A_{\sigma}\left(P_{\rho}+qA_{\rho}\right).$ (9)
When $A_{\mu}$ is independent of certain coordinates, say $x^{\hat{\mu}}$, Eq. (9) implies that the corresponding canonical momentum is an integral of motion. We reserve the symbol $Q_{\hat{\mu}}$ for these conserved momenta.
@@ III.1 The Low-Velocity Limit
Also, the action is manifestly invariant under general reparameterizations $\lambda\rightarrow\lambda^{\prime}(\lambda)$. We conclude that treating the temporal coordinate as a dynamical variable is not exclusive to relativistic theories. [^2] For an insightful discussion of reparametrization invariance and its relation time as a dynamical variable, see kieferQuantumGravity2004.
@@
$\dot{\mathcal{E}}$ $=\frac{q}{m}f\mathbf{E}(t,\mathbf{x})\!\cdot\!\mathbf{p}\,$, (16c)
$\dot{\mathbf{p}}$ $=\frac{q}{mc}f\left(\mathbf{E}(t,\mathbf{x})mc-\mathbf{B}(t,\mathbf{x})\!\times\!\mathbf{p}\right)\,$, (16d)
@@ III.2 The velocity sphere
which automatically satisfies $\mathbf{v}_{0}^{2}+\mu_{0}^{2}=1$. Therefore, a physically admissible initial state is fully determined by the set $\{\mathcal{E}_{0},\varphi_{0},\psi_{0}\}$, together with the initial position $\mathbf{x}_{0}\equiv\mathbf{x}(0)$, regardless of whether the particle is massive or massless.
@@ IV.1 Constant electromagnetic field
The last five equations define a closed dynamical subsystem in the variables $\{Z,\mathbf{Y}\}$. This structure suggests a solution strategy: first solve this reduced system, and then reconstruct the full trajectory $\{t(\lambda),\mathbf{x}(\lambda)\}$, using the integrals of motion
@@ IV.1.2 Numerical results
Dynamics of Charged Particles: Massive versus Massless.
@@
The representation of the orbit in terms of the $\xi$–coordinates (Fig. 1(c)) highlights the universal asymptotic behavior of charged particles, independent of their mass.
We now examine the dynamics in velocity space. As expected, the orbit of a massless particle remains strictly confined to the surface of the velocity sphere, whereas a massive particle can explore its interior (Fig. 1(b)). Qualitatively, the number of cycles described by the orbits is the same in both cases. For arbitrarily large positive (negative) values of $\lambda$, we have $\mathcal{E}\rightarrow\infty$, and both trajectories asymptotically approach the fixed point in the future (past), which can be computed by substituting the specific electromagnetic field parameters into Eqs. (21), yielding $\mathbf{v}^{\pm}=\left(0.1945,\pm 0.0161,\pm 0.9808\right)$.
@@
Numerical Comparison Between DTF and UTF.
@@ V Concluding Remarks
We have proposed a dynamical framework to describe the behavior of charged, massive and massless, test particles in the presence of external electromagnetic fields. This formulation is characterized by two main features: (i) the explicit implementation of the mass-shell relation as a constraint equation alongside the dynamical equations, and (ii) the inclusion of both time and energy as dynamical variables. The latter do not constitute a conjugate pair and should not be confused with the canonical variables appearing in the Hamiltonian formalism. Moreover, because $t$ is included among the dynamical variables, the resulting system of differential equations is always autonomous (in the sense that the evolution does not depend explicitly on the independent variable, $\lambda$). The incorporation of energy as a dynamical variable is essential in this approach, as it eliminates the need for the Lorentz factor, which becomes ill-defined in the limit of vanishing mass or numerically unstable for ultra-relativistic particles.
The dynamical-time framework provides a unified description of both the massive and massless sectors. The distinction between them arises solely from the choice of initial data, which must satisfy the mass-shell constraint that specifies the particle’s mass.
It should be emphasized that the evolution of the time coordinate does not represent a physical effect; rather, it reflects the underlying symmetry under general reparametrizations—a direct consequence of the mass-shell constraint being a first-class [henneauxQuantizationGaugeSystems1992].
@@
We have also examined the dynamics in velocity space, which offers a clear geometrical representation of several key aspects of charged-particle motion. The velocity sphere was defined as the two-dimensional manifold containing the trajectories of massless particles, while its interior is accessible only to massive ones. This construction enables both a qualitative and quantitative characterization of asymptotic behaviors, such as the tendency to achieve the speed of light for massive particles, as well as the existence of periodic and quasiperiodic regimes. [^4] See friedmanRelativisticAccelerationCharged2005 for a related approach based on the so-called s-velocity. Furthermore, with an appropriate parametrization of velocity space, one can readily specify sets of initial conditions consistent with the mass-shell constraint.
@@
Promoting the time coordinate to the status of a dynamical variable arises from a specific choice of parametrization, the so-called dynamical-time gauge. A notable feature of this gauge is that the subset of equations governing momentum and energy remains linear, albeit with coefficients locally depending on the spatial and temporal coordinates. Whether alternative gauges might or not provide analytical or numerical advantages remains an open question for future work.
From a numerical standpoint, the second part of this study highlights distinctive properties of the dynamical-time framework (DTF) in comparison to the conventional uniform-time framework (UTF). Both schemes enable accurate simulations of charged-particle dynamics, independent of particle mass. In particular, the mass-shell constraint and, where applicable, the integrals of motion are well preserved throughout the evolution. These findings were verified for both uniform and time-dependent electromagnetic field configurations.
Perhaps the most distinctive feature of the dynamical-time scheme—as suggested by our numerical results and valid specifically when an electric field directly drives the evolution of the time coordinate—is that the effective time coordinate can attain values far exceeding those reached in a uniform-time evolution, while adaptively adjusting its pace according to the intrinsic properties of the dynamical system. This “adaptive clock rate” effectively mimics, in many respects, the behavior of a variable-step numerical integrator operating within a uniform-time framework, which constitutes a capital result of the present study.
@@
The proposed framework, therefore, represents a viable and efficient alternative for numerical simulations. Despite the simplifications of our model, its ability to encompass the massless limit provides a consistent extension of charged-particle dynamics into the ultrarelativistic regime, offering new possibilities for both analytical and numerical exploration.
@@ Acknowledgements.
We thank Olivier Piguet for their insightful and valuable comments on this work. Z.O. is grateful to Lucy Aduviri for pointing out Ref. [LeBellac:1973unm]. GMRA thanks the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No 101034383.

### v2

@@ Abstract
We present a dynamical framework for modeling the motion of point-like charged particles, with or without mass, in general external electromagnetic fields. A key feature of this formulation is the treatment of time coordinate as a dynamical variable. The framework applies to the relativistic regime while consistently admitting a nonrelativistic limit. We also introduce a representation of particle trajectories in velocity space, which provides clear insight into the nature and asymptotic behavior of the dynamics. As an application, we compare the motion of massive and massless particles in a constant electromagnetic field and find that, for identical field configurations, their asymptotic behavior is independent of both mass and initial conditions. Finally, we explore the computational advantages of the dynamical-time formulation over the conventional uniform-time approach in two case studies: a uniform electromagnetic field, and an elliptically polarized wave propagating along a uniform magnetic field. In both scenarios, the proposed scheme exhibits improvements in accuracy and computational efficiency.
@@ I Introduction
A substantial body of literature addresses the classical dynamics of electrically charged particles, with or without spin and irrespective of their mass; a comprehensive review, however, lies beyond the scope of this work. A partial survey is provided in Ref. [moralesBehaviourChargedSpinning2018]. For complementary approaches not covered there, see Refs. [Plyushchay:1988ws, Plyushchay:1988wx, Nesterenko:1991av], as well as the extensive bibliography in Ref. [arreaga-garcia2014equations]. The dynamics of massive particles is well understood; for a complete analysis we refer to Ref. [Deriglazov2017]. Of particular relevance to the present study are the Lagrangian formulations in Refs. [brinkLocalSupersymmetrySpinning1976, brinkLagrangianFormulationClassical1977, balachandranClassicalDescriptionParticle1977], in which the kinetic term is decoupled from the mass, naturally encompassing the massless case.
In nonlinear dynamics, the behavior of charged particles in electromagnetic fields represents a fruitful area of research [akhiezerDynamicsHighenergyCharged1995, ramDynamicsChargedParticles2010]. For instance, massive charged particles in a dipolar magnetic field exhibit dynamical behavior ranging from periodic and quasiperiodic to chaotic and hyperchaotic in both the low-velocity [dealcantarabonfimChaoticHyperchaoticMotion2000] and relativistic regimes [shebalinStormerRegionsAxisymmetric2004, xiePeriodQuasiperiodChaos2020]. Extending such studies to massless particles could therefore provide valuable insights into ultrarelativistic dynamics. However, a primary difficulty in pursuing this extension is that, formally, the Lorentz factor diverges as the particle mass approaches zero, thereby challenging the applicability of standard approaches [petriFullyImplicitNumerical2017].
@@ II Dynamics of charged particles
In this section, we review the classical dynamics of charged particles, largely following Refs. [moralesBehaviourChargedSpinning2018, moralesQuantumChargedSpinning2019], whose conventions we adopt.
@@
$\tilde{S}[x,e]=-\int_{\mathcal{C}}\mathrm{d}\lambda\left(\frac{\dot{x}^{2}}{2e}+\frac{em^{2}}{2}+qA_{\mu}(x)\dot{x}^{\mu}\right),$ (2)
@@
In the action (2), the mass is decoupled from the kinetic term and appears as a free parameter multiplying the length of the particle’s worldline. As a consequence, the massless sector is realized in a straightforward manner. Moreover, one may even perform the substitution $m^{2}\rightarrow-m^{2}$ to describe tachyonic particles. [^1] The action (2) can be interpreted as a theory of (super)gravity in $0+1$ dimensions coupled to a multicomponent scalar field $x^{\mu}(\lambda)$ [brinkLocalSupersymmetrySpinning1976, brinkLagrangianFormulationClassical1977]. In this interpretation, $m^{2}$ plays the role of a cosmological constant, which may take negative values under the substitution $m^{2}\to-m^{2}$.
@@
The Hamiltonian formulation of this theory, in its most general form describing a charged spinning particle, was developed in Ref. [moralesQuantumChargedSpinning2019]. In the present case, where spin is neglected, the first-order action takes the form
$\tilde{S}[e,x,p]=\int\mathrm{d}\lambda\left(p_{\mu}\dot{x}^{\mu}-eH(x,p)\right),$ (5)
where $p_{\mu}\equiv\delta\tilde{S}/\delta\dot{x}^{\mu}$ are the canonical momenta computed from the action (2), and the Hamiltonian is given by
$H(x,p)=\frac{1}{2}\left(m^{2}-\eta^{\mu\nu}(p_{\mu}+qA_{\mu})(p_{\nu}+qA_{\nu})\right)\,.$ (6)
@@
For the canonically conjugate pair, satisfying $\{x^{\mu},p_{\nu}\}=\delta^{\mu}_{\nu}$, the Hamilton equations yield the following dynamical system:
@@
$\dot{p}_{\mu}$ $=\{p_{\mu},eH\}=qe\eta^{\nu\sigma}\partial_{\mu}A_{\nu}\left(p_{\sigma}+qA_{\sigma}\right).$ (9)
When $A_{\mu}$ is independent of certain coordinates, say $x^{\hat{\mu}}$, Eq. (9) implies that the corresponding canonical momentum is an integral of motion. We reserve the symbol $\mathcal{P}_{\hat{\mu}}$ for these conserved momenta.
@@ III.1 The Low-Velocity Limit
Also, the action is manifestly invariant under general reparameterizations $\lambda\rightarrow\lambda^{\prime}(\lambda)$. We conclude that treating the temporal coordinate as a dynamical variable is not exclusive to relativistic theories. [^2] For an insightful discussion of reparametrization invariance and its relation with time as a dynamical variable, see kieferQuantumGravity2004.
@@
$\dot{\mathcal{E}}$ $=f\frac{q}{m}\mathbf{E}(t,\mathbf{x})\!\cdot\!\mathbf{P}\,$, (16c)
$\dot{\mathbf{P}}$ $=f\frac{q}{mc}\left(\mathbf{E}(t,\mathbf{x})mc-\mathbf{B}(t,\mathbf{x})\!\times\!\mathbf{P}\right)\,$, (16d)
@@ III.2 The velocity sphere
which automatically satisfies $\mathbf{v}_{0}^{2}+\mu_{0}^{2}=1$. Therefore, a physically admissible initial state is fully determined by the set $(\mathcal{E}_{0},\varphi_{0},\psi_{0})$, together with the initial position $\mathbf{x}_{0}\equiv\mathbf{x}(0)$, regardless of whether the particle is massive or massless.
@@ IV.1 Constant electromagnetic field
The last five equations define a closed dynamical subsystem in the variables $(Z,\mathbf{Y})$. This structure suggests a solution strategy: first solve this reduced system, and then reconstruct the full trajectory $(t(\lambda),\mathbf{x}(\lambda))$, using the integrals of motion
@@ IV.1.2 Numerical results
Dynamics of Charged Particles: Massive versus Massless
@@
The representation of the orbit in terms of the $\xi$–coordinates (Fig. 1(b)) highlights the universal asymptotic behavior of charged particles, independent of their mass.
We now examine the dynamics in velocity space. As expected, the orbit of a massless particle remains strictly confined to the surface of the velocity sphere, whereas a massive particle can explore its interior (Fig. 1(c)). Qualitatively, the number of cycles described by the orbits is the same in both cases. For arbitrarily large positive (negative) values of $\lambda$, we have $\mathcal{E}\rightarrow\infty$, and both trajectories asymptotically approach the fixed point in the future (past), which can be computed by substituting the specific electromagnetic field parameters into Eqs. (21), yielding $\mathbf{v}^{\pm}=\left(0.1945,\pm 0.0161,\pm 0.9808\right)$.
@@
Numerical Comparison Between DTF and UTF
@@ V Concluding Remarks
We have proposed a dynamical framework to describe the motion of charged test particles, both massive and massless, in external electromagnetic fields. Our formulation relies on two key ingredients: a specific gauge choice for the einbein, which promotes time and energy to dynamical variables, and the free evolution of an eight-dimensional dynamical system, $(t,\mathbf{x};\mathcal{E},\mathbf{P})$, subject to initial data satisfying the mass-shell constraint. The variables introduced here should not be confused with the canonical variables of the Hamiltonian formalism. Not all degrees of freedom are physical or independent, since reparametrization invariance renders the evolution of $t$ physically irrelevant, and the mass-shell constraint requires only three of the four variables $(\mathcal{E},\mathbf{P})$ to define a well-posed initial-value problem. Consequently, the present framework provides a unified description of massive and massless sectors, distinguished solely by initial conditions encoding the particle’s mass.
This should be contrasted with constrained evolution, in which one typically adopts the uniform-time gauge, solves the mass-shell constraint to determine $\mathcal{E}=\sqrt{m^{2}+\mathbf{P}^{2}}$, and then evolves the remaining six variables using the evolution equations. It should be noted that this procedure, in principle, forbids negative-energy solutions, which are allowed classically for massless particles and underlie the so-called Klein tunneling phenomenon in quantum theory [moralesQuantumChargedSpinning2019].
Promoting time to a dynamical variable renders the system autonomous, while including energy eliminates the need for the Lorentz factor, a quantity that becomes ill-defined in the massless limit and numerically unstable in the ultrarelativistic regime.
From the initial-value formulation of the classical system, it follows that free, spinless, massless particles moving in the same direction and passing through the same spacetime point at $t(0)$ can be distinguished solely by their energy $\mathcal{E}_{0}$. This effective role of energy as inertia for massless particles, which was noted in Ref. [moralesBehaviourChargedSpinning2018], is confirmed by our numerical results. Invoking the correspondence principle, one may suggest that such particles are distinguished by their de Broglie wavelengths, $\lambda_{\mathrm{dB}}=hc/\mathcal{E}_{0}$. [^4] A general reformulation including spin degrees of freedom is required to further substantiate this correspondence. We thank Alexei Deriglazov for drawing our attention to this subtle point.
@@
We have also shown that the velocity-space representation provides an alternative perspective on charged-particle dynamics. The velocity sphere defines a submanifold that constrains the motion of massless particles, while its interior is accessible only to massive ones. This representation allows a clear characterization of asymptotic behaviors. Furthermore, with an angular parametrization of the velocity sphere, initial conditions consistent with the mass-shell constraint can be readily specified.
@@
The evolution of time, being proportional to the particle’s energy, arises from a specific parametrization choice, which we term the dynamical-time gauge. A notable feature of this gauge is that the subset of equations governing momentum and energy remains linear, although the coefficients depend locally on the spatial and temporal coordinates. Whether alternative gauge choices could provide analytical or numerical advantages remains an open question for future work.
From a numerical perspective, the dynamical-time framework (DTF) exhibits distinctive advantages over the conventional uniform-time framework (UTF). Both schemes provide accurate simulations, irrespective of the particle’s mass, with only minor violations of the mass-shell constraint and conserved quantities. Notably, as suggested by our numerical results, when electric fields directly drive the evolution of the time coordinate, the effective time variable adapts dynamically, reaching stages of evolution far exceeding those achievable in the uniform-time framework. This adaptive clock rate closely resembles the behavior of a variable-step integrator operating within a uniform-time framework. These results, verified for both uniform and time-dependent electromagnetic field configurations, constitute one of the main numerical findings of this work.
@@
Overall, the proposed framework represents a viable and efficient alternative for numerical simulations. Despite the simplifications of our model, its ability to encompass the massless limit provides a consistent extension of charged-particle dynamics into the ultrarelativistic regime, offering new opportunities for both analytical and numerical exploration.
@@ Acknowledgements.
We thank Alexei Deriglazov for extensive and insightful correspondence related to this work. We are also grateful to Olivier Piguet for a careful reading of the manuscript and for several valuable suggestions, and to Lucy Aduviry for bringing Ref. [LeBellac:1973unm] to our attention. GMRA acknowledges support from the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No. 101034383.